# Pretrained deep models outperform GBDTs in Learning-To-Rank under label scarcity

### v3

@@ 3 Design: Unsupervised pretraining for LTR
Pretraining method for LTR: Motivated by (a) SimCLR’s high complexity and (b) the efficacy of contrasting with hard negatives, we propose SimCLR-Rank, an LTR-specific alternative. Recall that the SimCLR loss for an item in (2) uses all the other items in the batch as negatives. Good pretraining strategies usually exploit the structure of the data, e.g. text sequences motivates masked token prediction (Devlin et al., 2018). Studying the query group structure of LTR data, we notice that high-quality hard negatives are freely available as the other items from the same query group. These items are retrieved and deemed potentially similarly relevant for this query by an upstream retrieval model, making these items harder negatives than other items in the batch. So, we propose SimCLR-Rank, which modifies SimCLR to contrast only with the other items in the same query group, as illustrated in Figure 2. Formally, we modify the per-augmentation loss $\ell^{(a)}_{q,i}$ as:
$\ell_{q,i}^{(a)}$ $=-\text{cos}(\boldsymbol{z}_{q,i}^{(a)},\boldsymbol{z}_{q,i}^{(\bar{a})})/\tau+\text{log}\sum_{i^{\prime}=1}^{L_{q}}\sum_{a^{\prime}=0}^{1}\big{[}$ (2)
$\mathbf{1}\{(i^{\prime},a^{\prime})\neq(i,a)\}\cdot\exp(\text{cos}(\boldsymbol{z}_{q,i}^{(a)},\boldsymbol{z}_{q,i^{\prime}}^{(a^{\prime})})/\tau)\big{]}$
@@ 3.1 Evaluation of unsupervised pretraining techniques
Dataset. We use MSLRWEB30K (or MSLR for brevity) (Qin & Liu, 2013), Yahoo Set1 (Chapelle & Chang, 2011), and Istella_S (Lucchese et al., 2016) (a smaller version of Istella (Dato et al., 2016)). In the tabular LTR literature, these three datasets are generally considered the standard; most recent papers evaluate (only) on these datasets (Qin et al., 2021; Ai et al., 2019; Pang et al., 2020; Yang et al., 2022; Yu, 2020). To simulate scarcity of labels, we let 0.2% of the training query groups in each dataset keep their labels, while the rest of the query groups have no labels. Dataset statistics are provided in Section A.2.1.
@@
Methodology. For the neural-based approaches, the base of the model is the tabular ResNet (Gorishniy et al., 2021). We pretrain on all the available query groups and finetune/train on only the labeled query groups. For the GBDT-based approaches, we use the implementation in lightgbm (Ke et al., 2017) and adopt a tuning strategy similar to Qin et al. (2021). We describe our experimental setup in detail in Section A.2.3. This includes hyperparameters, procedures for data augmentation, GBDT pseudo-labeling details, model training, and finetuning. For comparison to prior baselines on tabular data, our experimental setup is chosen to mimic the original setup in them.
@@
Results. Table 1 charts the test NDCG for each of the pretraining methods we study. SimCLR-Rank and SimSiam perform the best or the second best among these methods for most of the datasets. As expected, SimCLR-Rank is also 7-14x faster than SimCLR/SCARF/DACL+ (Table 7 in Section A.3.1). We also find that the semi-supervised GBDT baselines do not perform well—surprisingly, they tend to achieve lower NDCG than the GBDT baseline without semi-supervised learning. Given these advantages in terms of test NDCG and speed, we select SimCLR-Rank and SimSiam as the best overall pretraining methods for tabular LTR, and use them for our continued evaluation.
@@ 4 Empirical evaluation
Next we compare the best DL pretraining methods for tabular LTR (SimCLR-Rank and SimSiam) against non-pretrained DL and GBDT models on (1) the public datasets under different label scarcity patterns to simulate real-world scenarios [^1] Code is provided at https://anonymous.4open.science/r/ltr-pretrain-3C01/README.md., and (2) a private large-scale ranking dataset which is naturally label-scarce.
Table 2: Main result (Simulated implicit binary user feedback): When $X$of query groups are assigned a few binary labels, pretrained DL method outperform both GBDT and non-pretrained DL models. We generate labels with $\tau_{\text{target}}=4.5$ (Section 4.1.2). ${}^{\clubsuit}$ means that pretrained models are significantly better in NDCG than non-pretrained models, measured by a t-test with significance $p<0.05$.
MSLRWEB30K ($X=8.9\%$) Yahoo Set1 ($X=5.3\%$) Istella_S ($X=25.2\%$)
Method NDCG ($\uparrow$) Outlier-NDCG NDCG ($\uparrow$) Outlier-NDCG NDCG ($\uparrow$) Outlier-NDCG
GBDT 0.333 $\pm$ 0.000 0.288 $\pm$ 0.000 0.617 $\pm$ 0.000 0.540 $\pm$ 0.000 0.602 $\pm$ 0.000 0.678 $\pm$ 0.000
No Pretrain DL 0.323 $\pm$ 0.000 0.269 $\pm$ 0.029 0.610 $\pm$ 0.007 0.514 $\pm$ 0.005 0.608 $\pm$ 0.002 0.687 $\pm$ 0.015
Pretrained DL 0.356 $\pm$ ${}^{\clubsuit}$ 0.289 $\pm$ 0.012 0.622 $\pm$ ${}^{\clubsuit}$ 0.543 $\pm$ ${}^{\clubsuit}$ 0.613 $\pm$ ${}^{\clubsuit}$ 0.699 $\pm$ 0.003
@@ 4.1.1 Simulated crowdsourcing of labels
Methodology. We repeat the methodology from Section 3.1 for the GBDT and the SimCLR-Rank and SimSiam pretraining methods. Non-pretrained and pretrained DL rankers share the same architecture (see Section A.2.3 for more details). During hyperparameter tuning, we tune for NDCG and report the resulting test validation NDCG and test Outlier-NDCG. For each dataset and labeling fraction, we report the test metric of the best model amongst SimCLR-Rank and SimSiam under the name “pretrained DL” model.
Results. We compare pretrained rankers, non-pretrained DL rankers, and GBDTs on public datasets in Figure 1. The results for Outlier-NDCG are provided in Figure 5 in Section A.3.4. We find that pretrained rankers outperform non-pretrained methods (including GBDTs) on NDCG and Outlier-NDCG across all public datasets, up to a dataset-dependent fraction of QGs labeled, as shown in Figure 1 (for NDCG) and Figure 5 (for Outlier-NDCG). We also note that, GBDTs have better NDCG than non-pretrained DL models in most of these regimes. This suggests that, at least in some scenarios with a limited supply of labeled query groups, the self-supervised pretraining is the factor that allows deep learning models to outperform GBDTs.
@@ 4.1.2 Simulated implicit binary user feedback
Here we simulate a scenario when users of an LTR system provide binary implicit feedback for the items, like clicking of an advertisement on a webpage. This kind of feedback is usually infrequent in an LTR system, as it is the result of a user choosing to spend more time or resources on an item in a presented list. Therefore, most query groups will not have any labels, and even labeled query groups will only have a few positively labeled items. On the other hand, such labels are often cheaper to collect than crowdsourced labels, so they are used in many industry use cases.
@@
Methodology. We reuse the methodology of Section 4.1.1, except we use a linear scoring head for pretrained models to improve stability and performance (see Section A.2.3).
Results. In Table 2, we present the test NDCG and Outlier-NDCG metrics of the models under this setting. We see that pretrained rankers significantly outperform GBDTs and non-pretrained DL rankers in this simulated binary user feedback setting with up to, even with 25.2% labeled query groups (Istella_S). Note that, by default in lightgbm library, GBDT training has no randomness. Notice that for MSLRWEB30K and Yahoo Set1 datsets, the second best model is the GBDT, and the non-pretrained DL ranker has a much worse Outlier-NDCG. This aligns with the recent observation that GBDT models perform better than DL models in supervised training with datasets containing outliers (McElfresh et al., 2023). However, here pretraining helps DL rankers close this gap and at times beat GBDTs in outlier performance. We also present results for $\tau_{\text{target}}\in\{4.25,5.1\}$ in Section A.3.5. There, we find that making the labels sparser ($\tau_{\text{target}}=5.1$) increases the relative improvement of pretrained DL rankers, while in a less sparse setting ($\tau_{\text{target}}=4.25$) GBDTs dominate. This strongly suggests that pretrained DL models outperform GBDTs when learning from relatively rare implicit binary user feedback, which is common in search and recommendations.
@@ 4.2 Large-scale real-world dataset results
In this section we test whether the learnings from the simulated label scarce settings above, especially Section 4.1.2, translate to a large-scale real-world LTR problem.
Dataset. Our proprietary dataset is derived from the online shopping logs of a large online retailer, numbering in the tens of millions of query groups. In this dataset, query groups consist of items presented to a shopper if they enter a search query. We assign each item a purchase label: label is 1 if the shopper purchased the item, or 0 if the shopper did not purchase the item. Hence, this is a real-world instance of the implicit binary user feedback setting. Further, only about 10% of the query groups have items with non-zero labels. See Section A.2.1 for more dataset details.
@@
Table 3: Main result (Real-world dastaset): Results on a large-scale proprietary online shopping dataset. We give percentage improvement of rankers pretrained by SimSiam and SimCLR-Rank over GBDTs. Unsupervised pre-training improves (1) overall (full dataset) performance (internally, 2% is a big improvement), and (2) outlier performance.
@@
Results. Our results are summarized in Table 3 and all numbers are given as relative percentage improvements ($\Delta\%$) over the GBDT ranker. First, we note that all the DL models outperform GBDT models in terms of NDCG. Amongst them, the pretrained model using SimSiam performs the best with a gain of over $5.5\%$. This is substantial, given that in large-scale industry datasets, even a 2% improvement in NDCG is considered significant. In terms of Outlier-NDCG, the story is very different. Here, we see that the non-pretrained DL model and SimCLR-Rank performs up to -7% worse than the GBDT. However, SimSiam pretraining allows the DL ranker to overcome this shortcoming and beat the outlier performance of the GBDT. This verifies our hypothesis from Section 4.1.2 that pretraining can help DL model to outperform GBDT on large-scale real-world LTR problems, especially with implicit user feedback. Lastly, we see a significant performance difference between SimSiam and SimCLR-Rank. We will investigate this using an ablation study with public datasets later in (Section 4.3).
4.3 Peculiarities of best pretraining strategies for LTR

### v4

@@ 3 Design: Unsupervised pretraining for LTR
Pretraining method for LTR: Motivated by (a) SimCLR’s high complexity and (b) the efficacy of contrasting with hard negatives, we propose SimCLR-Rank, an LTR-specific alternative. Recall that the SimCLR loss for an item in equation 2 uses all the other items in the batch as negatives. Good pretraining strategies usually exploit the structure of the data, e.g. text sequences motivates masked token prediction Devlin et al. (2018). Studying the query group structure of LTR data, we notice that high-quality hard negatives are freely available as the other items from the same query group. These items are retrieved and deemed potentially similarly relevant for this query by an upstream retrieval model, making these items harder negatives than other items in the batch. So, we propose SimCLR-Rank, which modifies SimCLR to contrast only with the other items in the same query group, as illustrated in Figure 2. Formally, we modify the per-augmentation loss $\ell^{(a)}_{q,i}$ as:
$\ell_{q,i}^{(a)}$ $=-\text{cos}(\bm{z}_{q,i}^{(a)},\bm{z}_{q,i}^{(\bar{a})})/\tau+\text{log}\sum_{i^{\prime}=1}^{L_{q}}\sum_{a^{\prime}=0}^{1}\big{[}\mathbf{1}\{(i^{\prime},a^{\prime})\neq(i,a)\}\cdot\exp(\text{cos}(\bm{z}_{q,i}^{(a)},\bm{z}_{q,i^{\prime}}^{(a^{\prime})})/\tau)\big{]}$ (3)
@@ 3.1 Evaluation of unsupervised pretraining techniques
Table 1: Pretraining methods: A comparison across unsupervised pretraining methods on NDCG, averaged over 3 trials. Here only 0.1% of the query groups have labels for finetuning. Bold numbers denote the best in a column, and underlined numbers are within the margin of error of the best. Given that SCARF/DACL+/SimCLR are too slow to scale to large LTR datasets (for example on the industrial-scale proprietary online shopping dataset), we find that SimCLR-Rank and SimSiam are the best pretraining methods for tabular LTR.
Method MSLR ($\uparrow$) Yahoo Set1 ($\uparrow$) Istella ($\uparrow$)
Supervised GBDT 0.2801 $\pm$ 0.0002 0.5083 $\pm$ 0.0197 0.5450 $\pm$ 0.0000
Semi-supervised GBDT 0.2839 $\pm$ 0.0004 0.5061 $\pm$ 0.0267 0.4656 $\pm$ 0.0310
SCARF (Bahri et al., 2021) 0.3807 $\pm$ 0.0016 0.5884 $\pm$ 0.0129 0.5542 $\pm$ 0.0024
DACL+ (Verma et al., 2021) 0.3833 $\pm$ 0.0037 0.5887 $\pm$ 0.0012 0.5626 $\pm$ 0.0024
VIME-self (Yoon et al., 2020) 0.3834 $\pm$ 0.0011 0.5839 $\pm$ 0.0058 0.5514 $\pm$ 0.0025
SubTab (Ucar et al., 2021) 0.3748 $\pm$ 0.0025 0.5814 $\pm$ 0.0062 0.5082 $\pm$ 0.0050
SimCLR (Chen et al., 2020) 0.3827 $\pm$ 0.0027 0.5837 $\pm$ 0.0093 0.5602 $\pm$ 0.0055
SimCLR-Sample 0.3498 $\pm$ 0.0541 0.5803 $\pm$ 0.0070 0.5479 $\pm$ 0.0016
SimCLR-Rank 0.3868 $\pm$ 0.0026 0.5843 $\pm$ 0.0062 0.5609 $\pm$ 0.0040
SimSiam (Chen & He, 2021) 0.3790 $\pm$ 0.0028 0.6100 $\pm$ 0.0072 0.5189 $\pm$ 0.0096
@@
Dataset. We use MSLRWEB30K (or MSLR for brevity) (Qin & Liu, 2013), Yahoo Set1 (or Set1 for brevity) (Chapelle & Chang, 2011), and Istella_S (Lucchese et al., 2016) (a smaller version of Istella (Dato et al., 2016)). In the tabular LTR literature, these three datasets are generally considered the standard; most recent papers evaluate (only) on these datasets (Qin et al., 2021; Ai et al., 2019; Pang et al., 2020; Yang et al., 2022; Yu, 2020). To simulate scarcity of labels, we let 0.1% of the training query groups in each dataset keep their labels, while the rest of the query groups have no labels. Dataset statistics are provided in Section A.2.1.
Methodology. For the neural pretraining approaches (which include SimCLR-Rank, SimSiam, SimCLR, SubTab, VIME-self, DACL+, and SCARF), the base of the model is the tabular ResNet [^1] We emphasize that we do not prescribe any specific deep learning architecture choice; the choice for the best DL model may depend on the specific LTR task. Our proposed workflow of pretraining deep models in LTR is agnostic to the specific choice of DL architecture. Prior work has proposed DL architectures designed for click-through rate (CTR)/LTR based on factorization machines (Rendle, 2010), such as DeepFM and DCNv2 (Guo et al., 2017; Wang et al., 2021), which we compare to in Section 4.1.2 and Section 4.1.2 against our pretrained models. Note that it is unclear currently how to do unsupervised pretraining using these CTR/LTR specific architectures at the moment. (check the footnote for our explanation on the architecture choice). We also evaluate “SimCLR-Sample”, a neural pretraining baseline we create to demonstrate the value of negatives from the same query group. SimCLR-Sample is like SimCLR-Rank, but at the beginning of each epoch the items are first randomly permuted and grouped into fake query groups. Then we apply SimCLR-Rank. Thus SimCLR-Sample has the same time/space complexity as SimCLR-Rank. For the neural pretraining approaches (which include SimCLR-Rank, SimSiam, SimCLR, SubTab, VIME-self, DACL+, and SCARF), the base of the model is the tabular ResNet [^2] We emphasize that we do not prescribe any specific deep learning architecture choice; the choice for the best DL model may depend on the specific LTR task. Our proposed workflow of pretraining deep models in LTR is agnostic to the specific choice of DL architecture. Prior work has proposed DL architectures designed for click-through rate (CTR)/LTR based on factorization machines (Rendle, 2010), such as DeepFM and DCNv2 (Guo et al., 2017; Wang et al., 2021), which we compare to in Section 4.1.2 and Section 4.1.2 against our pretrained models. Note that it is unclear currently how to do unsupervised pretraining using these CTR/LTR specific architectures at the moment. (check the footnote for our explanation on the architecture choice). We pretrain on all the available query groups and finetune/train on only the labeled query groups.
@@
For the GBDT-based approaches, we use the implementation in lightgbm Ke et al. (2017) and adopt a tuning strategy similar to Qin et al. (2021). We report “Supervised GBDT” as GBDT utilizing only the labeled data, and in “Semi-supervised GBDT” we report the best among pseudo-labeled (Pseudo-Label, 2013) and PCA-enriched (Duh & Kirchhoff, 2008) GBDTs as well as the combination of the two techniques. We describe our experimental setup in detail in Section A.2.3. This includes hyperparameters, procedures for data augmentation, GBDT pseudo-labeling/PCA details, model training, and finetuning. For comparison to prior baselines on tabular data, our experimental setup is chosen to mimic the original setup in them.
@@
Results. Table 1 charts the test NDCG for each of the pretraining methods we study. SimCLR-Rank and SimSiam perform the best or the second best among these methods for most of the datasets. In particular, SimCLR-Rank outperforms SimCLR-Sample, demonstrating the value and efficiency of negatives from the same query group. As expected, SimCLR-Rank is also 7-14x faster than SimCLR/SCARF/DACL+ (Table 6 in Section A.3.1). We also find that the semi-supervised GBDT baselines do not perform well—surprisingly, they tend to achieve lower NDCG than the supervised GBDT baseline without semi-supervised learning. Given these advantages in terms of test NDCG and speed, we select SimCLR-Rank and SimSiam as the best overall pretraining methods for tabular LTR, and use them for our continued evaluation.
@@ 4 Empirical evaluation
Next we compare the best DL pretraining methods for tabular LTR (SimCLR-Rank and SimSiam) against GBDT and non-pretrained DL models on (1) the public datasets under different label scarcity patterns to simulate real-world scenarios [^3] Code is provided at https://anonymous.4open.science/r/ltr-pretrain-3C01/README.md., and (2) a private large-scale ranking dataset which is naturally label-scarce.
@@ 4.1.1 Simulated crowdsourcing of labels
Methodology. We repeat the methodology from Section 3.1 for the GBDT and the SimCLR-Rank and SimSiam pretraining methods. For each dataset and labeling fraction, we report the test metric of the best GBDT ranker amongst semi-supervised GBDTs and supervised GBDTs under “GBDT”. Similarly, under “no pretrain DL” we report the best supervised DL model amongst tabular ResNet (Gorishniy et al., 2021), DeepFM (Guo et al., 2017), and DCNv2 (Wang et al., 2021) under the name “no pretrain DL”. Finally, we report the best pretrained ranker amongst SimCLR-Rank and SimSiam under the name “pretrained DL”. Pretrained DL rankers use tabular ResNet (see Section A.2.3 for more details). During hyperparameter tuning, we tune for NDCG and report the resulting test NDCG and test Outlier-NDCG.
Results. We compare pretrained rankers, non-pretrained DL rankers, and GBDTs on public datasets in Figure 1. The results for Outlier-NDCG are provided in Figure 8 in Section A.3.4. We find that pretrained rankers outperform non-pretrained methods (including GBDTs) on NDCG and Outlier-NDCG across all public datasets, up to a dataset-dependent fraction of QGs labeled, as shown in Figure 1 (for NDCG) and Figure 8 (for Outlier-NDCG). We also note that GBDTs have better NDCG than non-pretrained DL models in most of these regimes. This shows that there exist scenarios with a limited supply of labeled query groups, where self-supervised pretraining is the factor that allows deep learning models to outperform GBDTs. We also provide a detailed a comparison between various semi-supervised and supervised GBDT models in Section A.4
@@ 4.1.2 Simulated implicit binary user feedback
Here we simulate a scenario when users of an LTR system provide binary implicit feedback for the items, like the clicking of an advertisement on a webpage. This kind of feedback is usually infrequent in an LTR system, as it is the result of a user choosing to spend more time or resources on an item in a presented list. Therefore, most query groups will not have any labels, and even labeled query groups will only have a few positively labeled items. On the other hand, such labels are often cheaper to collect than crowdsourced labels, so they are used in many industry use cases.
@@
Table 2: Main result (Simulated implicit binary user feedback): When $X$(MSLRWEB30K=8.9%, Yahoo Set1=5.3%, and in Istella_S=25.2%) of query groups are assigned a few binary labels, pretrained DL method outperform both GBDT and non-pretrained DL models. We generate labels with $\tau_{\text{target}}=4.5$ (Section 4.1.2). means that pretrained models are significantly better in NDCG than non-pretrained models, measured by a t-test with significance $p<0.05$.
Method MSLRWEB30K ($\uparrow$) Yahoo Set1 ($\uparrow$) Istella ($\uparrow$)
NDCG ($\uparrow$)
GBDT 0.346 $\pm$ 0.001 0.617 $\pm$ 0.000 0.602 $\pm$ 0.000
No Pretrain DL 0.333 $\pm$ 0.001 0.621 $\pm$ 0.000 0.608 $\pm$ 0.002
Pretrained DL 0.356 $\pm$ 0.004 0.622 $\pm$ 0.001 0.613 $\pm$ 0.002
Outlier-NDCG ($\uparrow$)
GBDT 0.288 $\pm$ 0.000 0.540 $\pm$ 0.000 0.678 $\pm$ 0.000
No Pretrain DL 0.238 $\pm$ 0.026 0.536 $\pm$ 0.003 0.687 $\pm$ 0.015
Pretrained DL 0.289 $\pm$ 0.012 0.543 $\pm$ 0.002 0.699 $\pm$ 0.003
Methodology. We reuse the methodology of Section 4.1.1, except we use a linear scoring head for pretrained models to improve stability and performance (see Section A.2.3). Similar to Section 4.1.1, we report the best of multiple models for each method and dataset (see Section 4.1.1 for details).
Results. In Table 2, we present the test NDCG and Outlier-NDCG metrics of the models under this setting. We see that pretrained rankers significantly outperform GBDTs and non-pretrained DL rankers in this simulated binary user feedback setting with up to, even with 25.2% labeled query groups (Istella_S). Notice that for MSLRWEB30K and Yahoo Set1 datasets, the second best model is the GBDT, and the non-pretrained DL ranker has a much worse Outlier-NDCG. This aligns with the recent observation that GBDT models perform better than DL models in supervised training with datasets containing irregularities McElfresh et al. (2023). However, here pretraining helps DL rankers close this gap and at times beat GBDTs in outlier performance. We also present results for $\tau_{\text{target}}\in\{4.25,5.1\}$ in Section A.3.5. There, we find that making the labels sparser ($\tau_{\text{target}}=5.1$) increases the relative improvement of pretrained DL rankers, while in a less sparse setting ($\tau_{\text{target}}=4.25$) GBDTs dominate. This strongly suggests that pretrained DL models outperform GBDTs when learning from relatively rare implicit binary user feedback, which is common in search and recommendation systems.
@@ 4.2 Large-scale real-world dataset results
In this section we test whether the results from the simulated label scarce settings above, especially Section 4.1.2, translate to a large-scale real-world LTR problem.
Dataset. Our proprietary dataset is derived from the online shopping logs of a large online retailer, numbering in the tens of millions of query groups. In this dataset, query groups consist of items presented to a shopper if they enter a search query. We assign each item a purchase label: label is 1 if the shopper purchased the item, or 0 if the shopper did not purchase the item. Hence, this is a real-world instance of the implicit binary user feedback setting simulated in Section 4.1.2. Further, only about 10% of the query groups have items with non-zero labels. See Section A.2.1 for more dataset details.
@@
Table 3: Main result (Real-world dataset): Results on a large-scale proprietary online shopping dataset. We give percentage improvement of rankers pretrained by SimSiam and SimCLR-Rank over GBDTs. Unsupervised pre-training improves (1) overall (full dataset) performance (internally, 2% is a big improvement), and (2) outlier performance.
@@
Results. Our results are summarized in Table 3 and all numbers are given as relative percentage improvements ($\Delta\%$) over the GBDT ranker. First, we note that all the DL models outperform GBDT models in terms of NDCG. Amongst them, the pretrained model using SimSiam performs the best with a gain of over $5.5\%$. This is substantial, given that in large-scale industry datasets, even a 2% improvement in NDCG is considered significant. In terms of Outlier-NDCG, the story is very different. Here, we see that the non-pretrained DL model and SimCLR-Rank performs up to -7% worse than the GBDT. However, SimSiam pretraining allows the DL ranker to overcome this shortcoming and beat the outlier performance of the GBDT. This verifies our hypothesis from Section 4.1.2 that pretraining can help DL models outperform GBDTs on large-scale real-world LTR problems, especially with implicit user feedback. Lastly, we see a significant performance difference between SimSiam and SimCLR-Rank. We will investigate this using an ablation study with public datasets later in (Section 4.3).
4.3 Ablations: Choices for pretrained models in LTR